OpenStudy (anonymous):

A tree casts a shadow that is 150 feet long. if the angle of elevation from the tip of the shadow to the top of the tree is 30 degrees, how tall is the tree? i know the answer is 87 feet. but why?

OpenStudy (anonymous):

 let the height of the tree be x feet. tan 30 = x/150 1/sqrt(3) = x/150 x=150/1.73 x=87 So, the height of the tree is 87 feet.

OpenStudy (anonymous):

Thank you, i got it know:)

OpenStudy (anonymous):

ur welcome :)